OpenStudy (anonymous):

Find the square. (2x-6)^2 A.4x^2-24x+36 B. 4x^2-8x+36 C. 4x^2-36 D. 4x^2-12x+36

mathslover (mathslover):

\[\large{(2x-6)^2=\color{blue}{(2x-6)(2x-6)}}\]

OpenStudy (anonymous):

( 2 x - 6 ) ( 2 x - 6 ) 4x^2 - 12 x - 12 x + 36 4 x^2 - 24 x + 36

mathslover (mathslover):

we have the formula for this: \[\large{\color{red}{(a-b)^2=a^2-2ab+b^2}}\]

mathslover (mathslover):

now put the given values of a = 2x and b = 6 in the above formula what do you get @jess.white ?

OpenStudy (anonymous):

that the answer no need to stress ur self

OpenStudy (anonymous):

4 x^2 - 24 x + 36

OpenStudy (anonymous):

lol @abayomi12

OpenStudy (anonymous):

yeah

mathslover (mathslover):

@abayomi12 : Please don't forget OPENSTUDY CODE-OF-CONDUCT it is not only this time that you're answering directly, you everytime do this work that is against code-of-conduct. for more info: http://openstudy.com/code-of-conduct but colm is here , he will handle that

OpenStudy (anonymous):

NO he doesnt he actually helps me he teaches me the steps on how to do the problems he is really a great teacher.

OpenStudy (anonymous):

hey guy i try to help people that don't know not all the math i can solve some u know some i don't ok?